Sodium reacts with chlorine gas according to the reaction: 2 Na(s) + Cl (9) ► 2 NaCl(s)

What volume of Cl2 gas, measured at 684 torr and 38 ∘C, is required to form 26 g of NaCl?

Express your answer using two significant figures.
Answer #1

Number of moles of NaCl = 26 g / 58.44 g/mol = 0.445 mole

From the balanced equation we can say that

2 mole of NaCl requires 1 mole of Cl2 so

0.445 mole of NaCl will require

= 0.445 mole of NaCl *(1 mole of Cl2 / 2 mole of NaCl)

= 0.223 mole of Cl2

PV = nRT

where, P = pressure = 684 torr = 0.900 atm

V = ?

n = number of moles

R = Gas constant

T = temperature = 38 + 273 = 311 K

0.900 * V = 0.223 * 0.0821 * 311

0.900 * V = 5.69

V = 5.69 / 0.900 = 6.3 L

Therefore, the volume of Cl2 required would be 6.3 L